OpenStudy (anonymous):

how do i do this problem find the exact solution of the expression sec^2pi/3-8

OpenStudy (anonymous):

\[\sec^2(\frac{\pi}{3})-8\]?

OpenStudy (anonymous):

yes exactly

OpenStudy (anonymous):

first find \(\cos(\frac{\pi}{3})\) do you know it ?

OpenStudy (anonymous):

but the directions say i would be able to do it without a calculator but i got .99983

OpenStudy (anonymous):

ooops

OpenStudy (anonymous):

take your calculator out of degree mode, put it in radian mode then it will give you the correct answer

OpenStudy (anonymous):

but in any case it would not hurt to know that \[\cos(\frac{\pi}{3})=\frac{1}{2}\]

OpenStudy (anonymous):

you can see it from the unit circle do you have a good unit circle cheat sheet?

OpenStudy (anonymous):

you do now in any case secant is the reciprocal of cosine the reciprocal of \(\frac{1}{2}\) is \(2\) square it, subtract 8 and you are done

OpenStudy (anonymous):

so -4?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you !!!!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

will you help me with another one ?

OpenStudy (anonymous):

sure